OpenStudy (anonymous):

A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. Part 1: What will the height of the ball be after 9 bounces? (4 points)

OpenStudy (anonymous):

~ 39 or 40

OpenStudy (anonymous):

Ok but how did you get that? I've tried it with a formula and the long way but i got 256,000 with the formula and 3.90625 doing it the long way.

OpenStudy (anonymous):

You need to use the geometric sequence. The common ratio r=0.5. The formula for the n-th term of the geometric sequence is:\[a_n = a_1 r^{n-1}\]Your a1=1000, and your n term is 9 because 9 bounces, so you replace that.\[a_9 = 1000*0.5^{9-1}=1000*0.5^8=3.90625\]

OpenStudy (anonymous):

Oh ok see when I did that i put 2 instead of .5. So 3.90625 is right?

OpenStudy (anonymous):

Yes. After 9 bounces you will get approximately 4 feet up.

OpenStudy (anonymous):

Ok thanks :)